OpenStudy (anonymous):

how do u do this problem 1178/166.97???

OpenStudy (anonymous):

very carefully

OpenStudy (anonymous):

can you manipulate the division so you are dividing by a whole number and not change the value of that fraction?

OpenStudy (anonymous):

what i dont get what yew are saying

OpenStudy (anonymous):

so...what you are doing is trying to divide correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so..to make things easier...you want to divide by a whole number...dividing by a decimal like 166.97 is tough So my question to you was...can you change that division, so that you have a whole number you divide by and not change the value of the problem

OpenStudy (anonymous):

soo like 1178/166,970

OpenStudy (anonymous):

okay nice!...you are getting the idea...but your numbers a bit off.

OpenStudy (anonymous):

so you changed the divisor by a multiple of 1000...what i mean is you multiplied the divisor (166.97) by 1000 to get 166970 but if you do that...you must also multipy the divident (1178) by 1000 or you change the value

OpenStudy (anonymous):

ok im in 6th grade and yew are talking on a 10th grade level can yew break it down please :)

OpenStudy (anonymous):

okay...one second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

this is what you are doing right?

OpenStudy (anonymous):

ya i think

OpenStudy (anonymous):

what i am suggesting is that you do something...so that the bottom of this fraction is a whole number...like this

OpenStudy (anonymous):

that's what you thought of when you gave me your last answer

OpenStudy (anonymous):

ok thats wat im doing

OpenStudy (anonymous):

okay good...BUT...you cannot just do that...because if you multipy by 1/1000 like that...then you are changing the value of your fraction

OpenStudy (anonymous):

so what do you think we can do...so that you don't change the value of your fraction...do something about that 1/1000

OpenStudy (anonymous):

try viewing my question

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

so mastery...do you have some clever way...to change that 1/1000 so that it won't mess with the value of your fraction?

OpenStudy (anonymous):

call me taylor i didt mean to put mastery and wat

OpenStudy (anonymous):

k, taylor...did they tell you in class that when you multiply anything by 1 then it does not change the value?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

so for example...: 5*1 = 5; 13*1 = 13; 1512*1 = 1512

OpenStudy (anonymous):

so anytime you multiply something by 1 ...you never change the value...that includes fractions

OpenStudy (anonymous):

i dont know anything about that

OpenStudy (anonymous):

ooooooo i think i get it

OpenStudy (anonymous):

so for example: 1/2 * 1 = 1/2 3/5 *1 =3/5 37383 / 37382 *1 = 37383/37382 and so on

OpenStudy (anonymous):

just any fraction times 1 is still the same fraction...

OpenStudy (anonymous):

okay the next thing you need to know is that if you have a fraction...that has the same number on top as on the bottom...then it's always equal to 1

OpenStudy (anonymous):

do you believe that?

OpenStudy (anonymous):

like... 3/3 = 1 or 5/5 =1 78/78 =1 1000/1000 = 1

OpenStudy (anonymous):

im really confused

OpenStudy (anonymous):

well, let's do one step at a time...do you believe my last statement...as long as the top number is equal to the bottom number it is equal to 1

OpenStudy (anonymous):

like this:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what im trying to say is that what you did earlier:

OpenStudy (anonymous):

is almost correct...but you cannot multiply by just 1/1000 you have to do it by 1000/1000...like this:

OpenStudy (anonymous):

because this way...you multiplied by something that is equal to 1 and that won't change the value of your fraction, but now you have a whole number on the bottom...and you can simplify or divide

OpenStudy (anonymous):

is there an easier way to do this

OpenStudy (anonymous):

not really the only thing you can do to make the numbers look a bit smaller is to use 100/100 instead of 1000/1000...cause 100/100 will do the trick...you went a bit overboard with multiplying by 1000/1000...but it gets the job done: so you can do this:

OpenStudy (anonymous):

see how doing it this way also gives you a whole number on the bottom...now you can simplify or divide as well

OpenStudy (anonymous):

i dont get what yew are saying!!

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

the hole thing'

OpenStudy (anonymous):

sorry you are gonna have to be more specific

OpenStudy (anonymous):

eveything u said so far

OpenStudy (anonymous):

please go back to the beginning of this thread and re-read what i wrote...hopefully you will have a more specific question once you do that

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

what was the last line that made sense to you?

OpenStudy (anonymous):

never mind i got it!!!!!!!!! thanx sooo much

OpenStudy (anonymous):

cool, you are welcome! I'm glad you didn't give up. Good luck

OpenStudy (anonymous):

thanx!